OpenStudy (anonymous):

Please help The perimeter of a rectangle is 96 ft. The ratio of the length to the width is 7: 5. What are the dimensions of the rectangle? A. 2x + 2x = 96 B. 7x + 5x = 96 C. 7x + 7x + 5x + 5x = 96 D. 7x/5x (fraction) = 96

OpenStudy (anonymous):

Second question Which is the best equation to use to solve this problem? The perimeter of a rectangle is 96 ft. The ratio of the length to the width is 7: 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (jack1):

that makes more sense ;D

OpenStudy (jack1):



OpenStudy (anonymous):

so for this first question it's C?

OpenStudy (jack1):

The ratio of the length to the width is 7: 5. so 7L = 5W so it's wider than it is long so if 7L = 5W than L = 5W/ 7

OpenStudy (jack1):

sub that into the perimeter equation, solve for W

OpenStudy (anonymous):

for the second question is it D?

OpenStudy (anonymous):

or is it perhaps B?

OpenStudy (jack1):

...? I'm not sure, sorry, I can give u the maths to work it out tho P = 2L + 2W 96 = 2L + 2W 96 = 2(5W/7) + 2W 96 = 10W / 7 + 2W 96 x 7 = 10W + 2W x7 672 = 10W + 14 W ... are you right to work out width from here?

OpenStudy (anonymous):

yes ^.^

OpenStudy (jack1):

sweet